Circular No. 6160

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995K IN ANONYMOUS GALAXY The High-Z Supernovae Search Team [B. Schmidt, Mount Stromlo and

Siding Spring Observatories; M. Phillips, M. Hamuy, R. Aviles, N. Suntzeff, and R. Schommer, Cerro Tololo Interamerican Observatory (CTIO); B. Leibundgut and J. Spyromilio, European Southern Observatory (ESO); B. Kirshner, A. Riess, P. Challis, and P. Garnavich, Harvard-Smithsonian Center for Astrophysics; and C. Stubbs and C. Hogan, University of Washington] report the discovery of a high-redshift supernova located 0′′.73 west and 0′′.82 north of a galaxy at α = 10^{h}50^{m}47^{s}.00, δ = −9^{o}15′07′′.4 (equinox 2000.0); SN 1995K is offset 11′′.4 east and 31′′.6 north from a nearby star (GSC 5499.0261; α = 10^{h}50^{m}46^{s}.2, δ = −9^{o}15′39′′). A finding chart can be obtained by e-mailing brian@merlin.anu.edu.au. SN 1995K was discovered (at mag R ∼ 22.5) on a frame taken with the CTIO 4-m telescope on Mar. 30.1 UT and is mariginally detected on a frame taken in poor conditions on Mar. 25. The supernova is not present on a deep image obtained on Mar. 7. Spectra obtained at the ESO New Technology Telescope on Apr. 3 suggest that the object is a type-Ia event slightly past maximum. The redshift of the host galaxy, derived from narrow Hα + [N II] emission, is z = 0.478.

AL COMAE BERENICES T. Augusteijn, European Southern Observatory, writes: “The dwarf

nova AL Com, currently in outburst (IAUC 6155), was observed spectroscopically (range 394–789 nm, resolution 0.9 nm) with the Danish 1.54-m telescope (+ DFOSC) at La Silla on Apr. 9.100 UT. A preliminary reduction shows shallow, wide absorption lines of Hα–ϵ and He I 447.1-nm (blended with Hγ, and possibly He I 438.8-nm), very similar to other dwarf novae in the early stages of a superoutburst. The total width of the lines at the continuum level ranges from ∼ 10000 km/s for Hα to ∼ 5500 km/s for Hδ. The equivalent widths of the lines range from ∼ 1.0 nm for Hα to 0.4 nm for Hδ. From a comparison with an observation of a flux standard, I find V ∼ 12.0 and B − V ∼ 0.0.”

Further visual magnitude estimates (cf. IAUC 6159): Apr. 11.864 UT, 13.0 (P. Schmeer, Bischmisheim, Germany); 12.191, 12.8 (W. G. Dillon, Missouri City, TX).

1995 April 12 Daniel W. E. Green